OpenStudy (anonymous):

Suppose that any given year, the value of a certain investment is increased by 15%. If the value is now $15,000, in how many years will the value be $21,000?

OpenStudy (anonymous):

use logarithms

OpenStudy (anonymous):

Anything and everything can be solved using logarithms

OpenStudy (anonymous):

\[21000 = 15000(1.15)^t\]

OpenStudy (anonymous):

And you can simplify that to be 1.4 = (1.15)^t....right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

solve for t

OpenStudy (anonymous):

Then take the log of both sides and solve for t

OpenStudy (anonymous):

\[1.4=(1.15)^t\] \[t=\frac{\ln(1.4)}{\ln(1.15)}\] in english "the log of the total divided by the log of the base"